# Towards Efficient 3D Object Detection in Bird’s-Eye-View Space for Autonomous Driving: A Convolutional-Only Approach

### v1

@@ I Introduction
3D object detection in Bird’s-Eye-View (BEV) space has gained considerable traction within the autonomous driving research community. As an alternative to LiDAR-based methods, generating pseudo-LiDAR points using surrounding cameras has emerged as a cost-effective and promising solution in the domain of autonomous driving. Consequently, numerous approaches [11, 15, 14, 10, 18, 6, 4, 8] have been proposed to incorporate perception tasks into the BEV space.
@@
However, existing methods are typically computationally demanding and heavily reliant on large-scale datasets. While these conditions can be met within laboratory settings, they present considerable obstacles to implementation in real-world, in-vehicle environments. The vision transformer (ViT) module [1], is the primary component responsible for substantial GPU memory consumption and matrix operations. Although the ViT architecture is widely utilized in BEV-based methods due to its capacity to capture global semantic information, it necessitates training on extensive datasets and requires a significantly longer training time than convolutional neural networks (CNNs) to facilitate the model’s understanding of the positional relationships between pixels. Despite the increased training costs, ViT offers only marginal improvements on various vision benchmarks compared to CNN-based models.
@@ II-A Backbone Models
Backbone models are the cornerstones of visual perception tasks. The inception of AlexNet [16] ignited a rapid progression of advancements in various vision tasks, with VGGNet [7] and ResNet [5] representing the first two significant milestones in backbone models for vision applications. VGGNet initially demonstrated the feasibility of enhancing vision task performance by increasing the depth of neural networks, while ResNet further expanded the depth of backbone models to 152 layers through the innovative use of skip connections. The Deformable Convolutional Network (DCN) [3] revealed that fixed filter sizes are not obligatory for learning visual features.
In the realm of efficient inference, EfficientNet [19] initially underscored the importance of deployment-friendly design, while RepVGG [13] recently established a burgeoning paradigm in inference-oriented models. Moreover, ElanNet [20], a vital embodiment of structural reparameterization, has showcased the unparalleled advantages of CNN models in deployment scenarios.
Beyond the field of computer vision, natural language processing (NLP) tasks have also reaped the benefits of advancements in backbone models. Transformers and their variants have dominated a considerable number of NLP tasks. Inspired by the widespread application of transformers in NLP, Dosovitskiy et al. proposed the Vision Transformer (ViT) [1] as a universal backbone model for computer vision. Drawing insights from dilated convolutional layers, Swin Transformer [17] represents another remarkable attempt to apply ViT to vision tasks.
@@
3D object detection, a fundamental task within the perception modules of autonomous driving systems, has experienced substantial advancements since the introduction of the BEV paradigm. Leveraging the precise semantic information provided by surrounding cameras, temporal input from previous frames, and depth information supervision from LiDAR inputs, camera-only 3D detection methods [10, 11, 15, 14, 6, 4, 8] are rapidly approaching the performance of their LiDAR-based counterparts.
DETR3D [23] pioneered the use of a multi-view image input paradigm to enhance the performance of 3D detection. Incorporating temporal information, BEVFormer [11] achieved a 10-point increase in mAP compared to DETR3D. BEVDet4D [15], which extended Lift-Splat-Shoot (LSS) [18] with an explicit BEV-Encoder, achieved similar results without requiring additional input from CAN bus information. BEVDepth [10], which incorporated depth supervision from LiDAR input, further validated the effectiveness of perception within the BEV space. Lift-Splat-Shoot [18], a classic yet efficient method for projecting images from a 2D-view to a 3D-view, initiated the utilization of surrounding camera input to enhance 3D perception performance. Each image is individually ”lifted” into a frustum of features for each camera, ”splat” all frustums into a rasterized bird’s-eye view grid, and finally ”shoot” different trajectories onto the cost map. With the emergence of BEV-based methods, it is anticipated that pure-vision-based approaches will soon match the performance of LiDAR-based methods.
@@ III-B Network Structure
Figure 4: Illustration of the Depth Module. We adopt the same design as BEVDepth [10] in depth estimation module, but add the augmentation matrix and extrinsic parameters together with the intrinsic parameters as input to the depth estimation network. The MLP layer is also being replaced by a convolutional network.
@@ III-C Backbone Model
In our quest to contrast the intricacies between ViT models and their convolutional counterparts, we strategically choose two representative models from each category. These include the naive Vision-Transformer (ViT) [1], Swin Transformer (SwinT) [17], the Resnet [5] and ElanNet [20]. Through this comparative study, we aim to unearth the best-performing model while adhering to our goals of improved simplicity and enhanced performance. Apart from the effort to compare the backbones, we simultaneously investigate the possible techniques to improve the model performance; we tried to mitigate the performance deterioration caused by the statistical offset of datasets from different sources. Specifically, pretraining on NuImage [2] has been adopted.
Refer to caption (a) Original Head Refer to caption (b) Merging Multi-Branch Parallel Conv Layers
@@ III-D View Projection
Following Lift-Splat-Shoot [18] and BEVDet [15], our feature projection module predicts depth probability for each pixel, calculating ground truth depth based on geometric similarity (Figure 3). Pixel $p$’s coordinates $(u,v)$ are transformed into feature map vector $f(u,v)$, representing depth probability at position $f(u,v)$ with vector $alpha\times f(u,v)$. Through light ray projection and the camera’s intrinsic and extrinsic parameters, image domain coordinates are converted to spatial domain. Final BEV feature computation happens via voxelization on a flat surface. The fixed 2D-to-3D projection matrix involves no learnable parameters. To accelerate inference, we compute this matrix during data processing and load it via a look-up table during training and testing.
@@ III-G Detection Head
Built upon the BEV feature, the detection head is based on CenterPoint [24]. We set the prediction target to include the position, scale, orientation and speed of the objects in the autonomous driving scenes. We adopt the same setup as CenterPoint during the training stage for a fair comparison against other algorithms. Formally we set the loss to be
@@
During the inference stage, as per RepVGG [13], we re-parameterize all the multi-branch convolutional layers and batch-norm layers into cascaded plain convolutional networks. As depicted in Figure 4(a), the detection head comprises several parallel convolutional neural networks. This structure can be simplified by merging the convolutional and batch-norm layers. As shown in Figure 4(c), a ResNet-like architecture is equivalent to a plain convolutional neural network without skip-connection or 1x1 convolution. The identity module can be directly added to the output feature map without any special operation. Meanwhile, the batch-norm layer can be combined with the convolutional layer by mathematically summing the mean value and standard variance of the batch input.
@@
FCOS3D [21] 1600x900 2008.2 1.7 29.5 37.2 0.806 0.268 0.511 1.315 0.170
DETR3D [23] 1600x900 1016.8 2.1 30.3 37.4 0.860 0.278 0.437 0.967 0.235
PGD [22] 1600x900 2223.0 1.4 33.5 40.9 0.732 0.263 0.423 1.285 0.172
BEVFormer [11] 1600x900 1303.5 1.7 41.6 47.6 0.673 0.274 0.372 0.394 0.198
@@
FCOS3D [21] 704x256 223.1 17.2 23.3 30.1 0.911 0.281 0.731 1.259 0.219
DETR3D [23] 704x256 195.5 19.7 23.9 31.0 0.972 0.292 0.625 1.073 0.304
PGD [22] 704x256 220.1 13.6 26.5 34.9 0.713 0.239 0.471 1.291 0.337
@@
BEVFormer [11] 704x256 217.2 19.9 32.8 39.5 0.661 0.259 0.357 1.593 0.197
BEVDet4D [14] 704x256 222.0 15.5 33.8 47.6 0.672 0.274 0.460 0.337 0.185
@@ IV-A Experiment Settings
Dataset and Evaluation Metrics Our model, BEVENet, is evaluated using the NuScenes benchmark dataset [2], encompassing 1,000 driving scenes captured via six cameras and a LiDAR sensor. This dataset annotates ten classes within a 51.2-meter ground plane for the 3D detection task. Performance evaluation leverages both official NuScenes metrics, namely, mean Average Precision (mAP), Average Translation Error (ATE), Average Scale Error (ASE), Average Orientation Error (AOE), Average Velocity Error (AVE), Average Attribute Error (AAE), and NuScenes Detection Scores (NDS), as well as efficiency-oriented metrics, namely, Frame Per Second (FPS) and GFlops. The former measures performance on NVIDIA A100 GPU with pre-processing and post-processing time excluded, while the latter uses the MMDetection3D [12] toolkit.
@@
Data Processing Our data processing adopts an approach similar to BEVDet [15], tailored for the specific needs of the NuScenes dataset which has an original resolution of $1600\times 900$. We rescale this to $704\times 256$ during preprocessing. Key transformations include random flipping, scaling, cropping, rotation, and a copy-paste mechanism to address any skewness in object distribution. These augmentation operations are mathematically converted into transformation matrices, combined with camera parameters that are flattened for dimension consistency. Class-Balanced-Grouping-and-Sampling (CBGS) [25], in concert with the copy-paste mechanism, is applied during training, following the methodology of CenterPoint [24]. In testing phase, images are scaled only but not cropped to align with the model’s input dimensions.
@@ IV-B Performance Benchmark
We select 11 SOTA methods on NuScenes leaderboard as our baselines: BEVFormer [11], BEVDet [15], BEVDet4D [14], BEVDepth [10], PETR [9], PGD [22], FCOS3D [21], DETR3D [23], CAPE [6], SoloFusion [4] and TiGBEV [8].
@@ V-B Backbone Module
In analyzing different backbone models, we scrutinized ViT [1], SwinTransformer [17], ResNet [5], and ELanNet [20], each with a similar parameter count for impartial comparison. Starting with ResNet50 as our baseline, with FPS at 27.4, mAP at 39.3 and NDS at 45.9 as per Table IIIIII(a), we found ViT underperforming in FPS at 17.9, Swin-T on par with ResNet50, and ELanNet superior in FPS at 30.2. Augmenting ELanNet with NuImage pretraining enhanced performance notably, increasing FPS to 30.3 while improving mAP and NDS to 42.0 and 50.1, respectively. Therefore, ELanNet, designed for inference efficiency, outpaced ViT, Swin-Transformer, and ResNet50 in this task.
@@ V-C View Projection and Depth Estimation Modules
Simultaneously, the depth estimation module, vital for post-projection depth optimization, was examined. Initially mirroring BEVDepth structure [10], Figure 6 revealed its substantial complexity contribution of 10.9%. Consequently, the MLP layer was replaced with a 2-layer residual block, a minimalist redesign elevating FPS by nearly one point and mAP by two points (Table IIIIII(c)). Moreover, we further enhanced it by integrating the image augmentation matrix. Despite the seemingly marginal impact, our experimental results confirmed the efficacy of this strategy in bolstering model performance.

### v2

# prof_yeo.bib

@@ I Introduction
3D object detection in Bird’s-Eye-View (BEV) space has gained considerable traction within the autonomous driving research community. As an alternative to LiDAR-based methods, generating pseudo-LiDAR points using surrounding cameras has emerged as a cost-effective and promising solution in the domain of autonomous driving. Consequently, numerous approaches [li2022bevformer, huang2021bevdet, huang2022bevdet4d, li2022bevdepth, philion2020lift, cape2023xiong, solofusion2022park, tigbev2022huang] have been proposed to incorporate perception tasks into the BEV space.
However, existing methods are typically computationally demanding and heavily reliant on large-scale datasets. While these conditions can be met within laboratory settings, they present considerable obstacles to implementation in real-world, in-vehicle environments. The vision transformer (ViT) module [dosovitskiy2020vit], is the primary component responsible for substantial GPU memory consumption and matrix operations. Although the ViT architecture is widely utilized in BEV-based methods due to its capacity to capture global semantic information, it necessitates training on extensive datasets and requires a significantly longer training time than convolutional neural networks (CNNs) to facilitate the model’s understanding of the positional relationships between pixels. Despite the increased training costs, ViT offers only marginal improvements on various vision benchmarks compared to CNN-based models.
@@ II-A Backbone Models
Backbone models are the cornerstones of visual perception tasks. The inception of AlexNet [alex2012alexnet] ignited a rapid progression of advancements in various vision tasks, with VGGNet [olga2015vgg] and ResNet [He2015] representing the first two significant milestones in backbone models for vision applications. VGGNet initially demonstrated the feasibility of enhancing vision task performance by increasing the depth of neural networks, while ResNet further expanded the depth of backbone models to 152 layers through the innovative use of skip connections. The Deformable Convolutional Network (DCN) [dai2017dcn] revealed that fixed filter sizes are not obligatory for learning visual features.
In the realm of efficient inference, EfficientNet [tan2019efficient] initially underscored the importance of deployment-friendly design, while RepVGG [ding2021repvgg] recently established a burgeoning paradigm in inference-oriented models. Moreover, ElanNet [wang2022yolov7], a vital embodiment of structural reparameterization, has showcased the unparalleled advantages of CNN models in deployment scenarios.
Beyond the field of computer vision, natural language processing (NLP) tasks have also reaped the benefits of advancements in backbone models. Transformers and their variants have dominated a considerable number of NLP tasks. Inspired by the widespread application of transformers in NLP, Dosovitskiy et al. proposed the Vision Transformer (ViT) [dosovitskiy2020vit] as a universal backbone model for computer vision. Drawing insights from dilated convolutional layers, Swin Transformer [liu2021Swin] represents another remarkable attempt to apply ViT to vision tasks.
@@
3D object detection, a fundamental task within the perception modules of autonomous driving systems, has experienced substantial advancements since the introduction of the BEV paradigm. Leveraging the precise semantic information provided by surrounding cameras, temporal input from previous frames, and depth information supervision from LiDAR inputs, camera-only 3D detection methods [li2022bevdepth, li2022bevformer, huang2021bevdet, huang2022bevdet4d, cape2023xiong, solofusion2022park, tigbev2022huang] are rapidly approaching the performance of their LiDAR-based counterparts.
DETR3D [detr3d] pioneered the use of a multi-view image input paradigm to enhance the performance of 3D detection. Incorporating temporal information, BEVFormer [li2022bevformer] achieved a 10-point increase in mAP compared to DETR3D. BEVDet4D [huang2021bevdet], which extended Lift-Splat-Shoot (LSS) [philion2020lift] with an explicit BEV-Encoder, achieved similar results without requiring additional input from CAN bus information. BEVDepth [li2022bevdepth], which incorporated depth supervision from LiDAR input, further validated the effectiveness of perception within the BEV space. Lift-Splat-Shoot [philion2020lift], a classic yet efficient method for projecting images from a 2D-view to a 3D-view, initiated the utilization of surrounding camera input to enhance 3D perception performance. Each image is individually ”lifted” into a frustum of features for each camera, ”splat” all frustums into a rasterized bird’s-eye view grid, and finally ”shoot” different trajectories onto the cost map. With the emergence of BEV-based methods, it is anticipated that pure-vision-based approaches will soon match the performance of LiDAR-based methods.
@@ III-B Network Structure
Figure 4: Illustration of the Depth Module. We adopt the same design as BEVDepth [li2022bevdepth] in depth estimation module, but add the augmentation matrix and extrinsic parameters together with the intrinsic parameters as input to the depth estimation network. The MLP layer is also being replaced by a convolutional network.
@@ III-C Backbone Model
In our quest to contrast the intricacies between ViT models and their convolutional counterparts, we strategically choose two representative models from each category. These include the naive Vision-Transformer (ViT) [dosovitskiy2020vit], Swin Transformer (SwinT) [liu2021Swin], the Resnet [He2015] and ElanNet [wang2022yolov7]. Through this comparative study, we aim to unearth the best-performing model while adhering to our goals of improved simplicity and enhanced performance. Apart from the effort to compare the backbones, we simultaneously investigate the possible techniques to improve the model performance; we tried to mitigate the performance deterioration caused by the statistical offset of datasets from different sources. Specifically, pretraining on NuImage [nuscenes2019] has been adopted.
(a) Original Head
(b) Merging Multi-Branch Parallel Conv Layers
@@ III-D View Projection
Following Lift-Splat-Shoot [philion2020lift] and BEVDet [huang2021bevdet], our feature projection module predicts depth probability for each pixel, calculating ground truth depth based on geometric similarity (Figure 3). Pixel $p$’s coordinates $(u,v)$ are transformed into feature map vector $f(u,v)$, representing depth probability at position $f(u,v)$ with vector $alpha\times f(u,v)$. Through light ray projection and the camera’s intrinsic and extrinsic parameters, image domain coordinates are converted to spatial domain. Final BEV feature computation happens via voxelization on a flat surface. The fixed 2D-to-3D projection matrix involves no learnable parameters. To accelerate inference, we compute this matrix during data processing and load it via a look-up table during training and testing.
@@ III-G Detection Head
Built upon the BEV feature, the detection head is based on CenterPoint [yin2021center]. We set the prediction target to include the position, scale, orientation and speed of the objects in the autonomous driving scenes. We adopt the same setup as CenterPoint during the training stage for a fair comparison against other algorithms. Formally we set the loss to be
@@
During the inference stage, as per RepVGG [ding2021repvgg], we re-parameterize all the multi-branch convolutional layers and batch-norm layers into cascaded plain convolutional networks. As depicted in Figure 5(a), the detection head comprises several parallel convolutional neural networks. This structure can be simplified by merging the convolutional and batch-norm layers. As shown in Figure 5(c), a ResNet-like architecture is equivalent to a plain convolutional neural network without skip-connection or 1x1 convolution. The identity module can be directly added to the output feature map without any special operation. Meanwhile, the batch-norm layer can be combined with the convolutional layer by mathematically summing the mean value and standard variance of the batch input.
@@
FCOS3D [wang2021fcos3d] 1600x900 2008.2 1.7 29.5 37.2 0.806 0.268 0.511 1.315 0.170
DETR3D [detr3d] 1600x900 1016.8 2.1 30.3 37.4 0.860 0.278 0.437 0.967 0.235
PGD [wang2021pgd] 1600x900 2223.0 1.4 33.5 40.9 0.732 0.263 0.423 1.285 0.172
BEVFormer [li2022bevformer] 1600x900 1303.5 1.7 41.6 47.6 0.673 0.274 0.372 0.394 0.198
@@
FCOS3D [wang2021fcos3d] 704x256 223.1 17.2 23.3 30.1 0.911 0.281 0.731 1.259 0.219
DETR3D [detr3d] 704x256 195.5 19.7 23.9 31.0 0.972 0.292 0.625 1.073 0.304
PGD [wang2021pgd] 704x256 220.1 13.6 26.5 34.9 0.713 0.239 0.471 1.291 0.337
@@
BEVFormer [li2022bevformer] 704x256 217.2 19.9 32.8 39.5 0.661 0.259 0.357 1.593 0.197
BEVDet4D [huang2022bevdet4d] 704x256 222.0 15.5 33.8 47.6 0.672 0.274 0.460 0.337 0.185
@@ IV-A Experiment Settings
Dataset and Evaluation Metrics Our model, BEVENet, is evaluated using the NuScenes benchmark dataset [nuscenes2019], encompassing 1,000 driving scenes captured via six cameras and a LiDAR sensor. This dataset annotates ten classes within a 51.2-meter ground plane for the 3D detection task. Performance evaluation leverages both official NuScenes metrics, namely, mean Average Precision (mAP), Average Translation Error (ATE), Average Scale Error (ASE), Average Orientation Error (AOE), Average Velocity Error (AVE), Average Attribute Error (AAE), and NuScenes Detection Scores (NDS), as well as efficiency-oriented metrics, namely, Frame Per Second (FPS) and GFlops. The former measures performance on NVIDIA A100 GPU with pre-processing and post-processing time excluded, while the latter uses the MMDetection3D [mmdet3d2020] toolkit.
@@
Data Processing Our data processing adopts an approach similar to BEVDet [huang2021bevdet], tailored for the specific needs of the NuScenes dataset which has an original resolution of $1600\times 900$. We rescale this to $704\times 256$ during preprocessing. Key transformations include random flipping, scaling, cropping, rotation, and a copy-paste mechanism to address any skewness in object distribution. These augmentation operations are mathematically converted into transformation matrices, combined with camera parameters that are flattened for dimension consistency. Class-Balanced-Grouping-and-Sampling (CBGS) [zhu2019class], in concert with the copy-paste mechanism, is applied during training, following the methodology of CenterPoint [yin2021center]. In testing phase, images are scaled only but not cropped to align with the model’s input dimensions.
@@ IV-B Performance Benchmark
We select 11 SOTA methods on NuScenes leaderboard as our baselines: BEVFormer [li2022bevformer], BEVDet [huang2021bevdet], BEVDet4D [huang2022bevdet4d], BEVDepth [li2022bevdepth], PETR [liu2022petr], PGD [wang2021pgd], FCOS3D [wang2021fcos3d], DETR3D [detr3d], CAPE [cape2023xiong], SoloFusion [solofusion2022park] and TiGBEV [tigbev2022huang].
@@ V-B Backbone Module
In analyzing different backbone models, we scrutinized ViT [dosovitskiy2020vit], SwinTransformer [liu2021Swin], ResNet [He2015], and ELanNet [wang2022yolov7], each with a similar parameter count for impartial comparison. Starting with ResNet50 as our baseline, with FPS at 27.4, mAP at 39.3 and NDS at 45.9 as per Table IIIIII(a), we found ViT underperforming in FPS at 17.9, Swin-T on par with ResNet50, and ELanNet superior in FPS at 30.2. Augmenting ELanNet with NuImage pretraining enhanced performance notably, increasing FPS to 30.3 while improving mAP and NDS to 42.0 and 50.1, respectively. Therefore, ELanNet, designed for inference efficiency, outpaced ViT, Swin-Transformer, and ResNet50 in this task.
@@ V-C View Projection and Depth Estimation Modules
Simultaneously, the depth estimation module, vital for post-projection depth optimization, was examined. Initially mirroring BEVDepth structure [li2022bevdepth], Figure 6 revealed its substantial complexity contribution of 10.9%. Consequently, the MLP layer was replaced with a 2-layer residual block, a minimalist redesign elevating FPS by nearly one point and mAP by two points (Table IIIIII(c)). Moreover, we further enhanced it by integrating the image augmentation matrix. Despite the seemingly marginal impact, our experimental results confirmed the efficacy of this strategy in bolstering model performance.